# SwiftSeg: Efficient Training-Free Open-Vocabulary Segmentation via Hierarchical Attention Refinement Method

### v2

@@ 4.0.1 Datasets and Metric
We evaluate the proposed SwiftSeg on standard benchmark datasets for the Open-Vocabulary Semantic Segmentation (OVSS) task: PASCAL VOC, PASCAL Context, and MS COCO Object, comprising 20, 59, and 80 object categories, respectively (excluding the background class). As SwiftSeg is a training-free method, evaluation is conducted directly on the validation sets of each dataset, which include 1,449 images from PASCAL VOC, 5,004 images from PASCAL Context, and 5,000 images from MS COCO Object. To assess segmentation performance and enable fair comparison with existing approaches, we employ the mean Intersection-over-Union (mIoU) metric, which is a widely used standard for evaluating semantic segmentation models that measure the overlap between predicted and ground-truth regions across classes.
@@ 4.0.2 Implementation Details
We employ 2-Rectified Flow, a distilled variant of Stable Diffusion 1.5 (SD 1.5), as the primary model for attention map extraction. All inferences of SD are performed at an image resolution of 512$\times$512 pixels. In the multi-resolution attention fusion module, we assign distinct weights to cross-attention and self-attention maps. Specifically, for cross-attention maps, we use the weight vector $\{0.15,0.7,0.15,0\}$ corresponding to resolution levels $\{8,16,32,64\}$, while self-attention maps, we assign weights of $\{0.1,0.1,0.5,0.3\}$. These weights are determined through both qualitative inspection of attention heatmaps and quantitative evaluation of segmentation performance at each resolution (see Table 5). Rather than relying on extensive hyperparameter tuning, we prioritize semantic relevance, assigning greater weights to layers that contribute more effectively to accurate segmentation. All experiments are implemented in PyTorch and conducted on a system equipped with an NVIDIA GeForce RTX A5000 GPU (32GB RAM) and an Intel(R) Core(TM) i9-10900X CPU. SwiftSeg is executed using mixed precision (16-bit floating point), requiring approximately 13.4 GB of GPU memory with a batch size of 2, and achieves an average inference time of 0.36 seconds per image (excluding candidate class extraction).
@@ 4.1.1 Quantitative results
Table 1 summarizes the performance of our proposed SwiftSeg method across three widely-used open-vocabulary semantic segmentation benchmarks: PASCAL VOC 2012, PASCAL Context, and COCO Object. SwiftSeg achieves mIoU scores of 64.0% on PASCAL VOC, 31.3% on PASCAL Context, and 36.2% on COCO Object, leading to the highest average mIoU among training-free diffusion-based approaches across all three datasets. While results from contrastive vision-language models are also reported for comprehensive comparison, our primary focus centers on a direct and fair comparison with recent training-free diffusion-based methods, including Diffusion Baseline [39], FreeSeg-Diff [19], DiffCut [38], DiffSeg [36], DiffSegmenter [20], and InvSeg [21]. The performance of Diffusion Baseline is referenced from the InvSeg study [21], while the results of Semantic DiffSeg, an adaptation of DiffSeg [36] for open-vocabulary settings, are taken from its corresponding publication [20]. SwiftSeg establishes state-of-the-art accuracy on PASCAL VOC 2012 and delivers the highest overall average performance across all datasets. On PASCAL Context and COCO Object, SwiftSeg ranks second, with results only marginally trailing behind the best-performing methods on each dataset.
@@
SwiftSeg (Our) RTX 4090 13.4G 0.36s 43.8
Figure 8: Open vocabulary semantic segmentation results of SwiftSeg and DiffSegmenter [20], DiffCut [38], ClearCLIP [5], CLIP-DIY [3], OVSegmentor [12]
Table 3: Ablation study on the contributions of each component in SwiftSeg, including self-attention refinement, cross-attention weighted aggregation, self-attention weighted aggregation, class prompt design, and test-time flipping. Removing any individual component leads to decreased performance.
SwiftSeg PASCAL VOC PASCAL Context COCO Object $\downarrow\Delta$
@@
w/o cross attention weighted aggregation 61.2 30.2 35.0 1.4
@@
Table 2 presents a quantitative evaluation of the operational efficiency among existing state-of-the-art diffusion-based segmentation methods, including DiffSegmentor, InvSeg, and our SwiftSeg. In terms of memory consumption, SwiftSeg demonstrates superior efficiency, utilizing only 13.4GB, significantly less than DiffSegmentor’s 16.1GB and InvSeg’s 32.4GB . Notably, regarding inference time, SwiftSeg achieves superior performance with just 0.36 seconds on an RTX 4090 GPU. This figure represents a substantial improvement over DiffSegmentor, whose inference time depends on the number of candidate classes M, and is particularly impressive when compared to InvSeg, which requires up to 7.9 seconds per image on an NVIDIA H100 GPU due to its complex prompt embedding optimization procedure. These results demonstrate that SwiftSeg offers high inference efficiency while significantly reducing the necessary computational resources.
In summary, SwiftSeg not only achieves state-of-the-art segmentation performance among training-free diffusion-based methods but also demonstrates superior inference efficiency and deployment simplicity, making it an efficient and scalable solution for real-world open-vocabulary semantic segmentation.
@@ 4.1.2 Qualitative results
Figure 8 provides a qualitative comparison of SwiftSeg against several methods, including DiffSegmenter, DiffCut, ClearCLIP, CLIP-DIY, and OVSegmentor, on two benchmark datasets: PASCAL VOC 2012 (top 7 rows) and PASCAL Context (bottom 6 rows). To ensure fairness, the PAMR post-processing step [46] was removed when showing DiffCut results. The results highlight several key observations. DiffSegmenter performs well in capturing object shapes on PASCAL VOC but struggles with contextual classes (e.g., tree, glass) in PASCAL Context. Additionally, its performance is highly sensitive to the quality of the candidate class list. DiffCut produces unstable segmentation masks, attributed to its reliance on an unsupervised grouping strategy and requiring FC-CLIP for semantic label association, making it suboptimal for open-vocabulary segmentation tasks. Among CLIP-based methods, CLIP-DIY often suffers from false positives and false negatives, while ClearCLIP partially mitigates this issue but still shows failure cases—particularly missing detections for classes like person. This indicates limited spatial localization capacity in these models. OVSegmentor exhibits strong geometric segmentation but lacks effective mechanisms to filter out irrelevant or negative classes, leading to noisy outputs and reduced precision.
In contrast, SwiftSeg consistently delivers high-quality segmentation masks, demonstrating both accurate geometric localization and precise category alignment, even in challenging scenarios. These results substantiate SwiftSeg’s advantage in open-vocabulary settings, offering a strong balance between semantic accuracy and spatial precision.
@@ 4.2.1 Effects of different components
Table  3 presents an ablation study that evaluates the contribution of components in SwiftSeg. Removing the cross-attention and self-attention weighted aggregation modules is equivalent to using only a single-resolution attention map with the best performance (as shown in Table  5), without applying multi-resolution aggregation-specifically, $16\times 16$ for cross-attention and $32\times 32\times 32\times 32$ for self-attention. In contrast, removing the class prompt $\mathcal{P}^{\prime}$ means that cross-attention maps are computed solely based on the text prompt $\mathcal{P}$, without explicitly guiding the model to focus on candidate classes. In the original configuration, the class candidate list is appended to the image caption to form the complete prompt $\mathcal{P}$, allowing the model to utilize the appended candidate list to generate more targeted cross-attention maps. Among all components, self-attention refinement proves to be the most critical, as its removal leads to the most significant average performance drop (-15.1%), underscoring its importance in enhancing the spatial precision of class-aware attention. Both cross-attention and self-attention weighted aggregation also play important roles, as their removal results in noticeable performance degradation. On the other hand, class prompt design and test-time flipping (TTF) have relatively more minor effects-0.6% and 0.3%, respectively-yet consistently contribute to overall segmentation quality.
These findings validate the overall design of SwiftSeg, where each component serves a complementary role in improving segmentation performance while maintaining computational efficiency.
@@
4.2.2 Different background threshold
@@
Table 6: Ablating different class candidate generation strategies.
PASCAL VOC VOC Context COCO Object
BIP 58.3 26.4 34.3
CLIP 63.0 23.9 34.8
TagCLIP 65.0 29.3 34.8
CLIP+BIP 62.8 28.8 36.7
TagCLIP+BIP 64.0 31.3 36.2
Ground Truth 69.7 45.1 38.8
4.2.3 Different class candidate generation strategies.
4.2.4 Effect of different attention scales
Table 5 presents a comprehensive study on the effect of different attention scales and fusion strategies in SwiftSeg. The results reveal several key findings. First, single-resolution attention maps, particularly those from the 16×16 resolution level, consistently yield higher segmentation performance across all datasets compared to other individual scales, suggesting that this resolution offers a favorable balance between semantic context and spatial detail. Second, when combining attention maps across multiple resolutions, both mean and weighted aggregation strategies improve overall performance, highlighting the complementary nature of multi-scale attention. Notably, weighted fusion outperforms mean fusion, as it enables the model to emphasize semantically rich and informative layers (e.g., 16×16 and 32×32) while down-weighting noisier or less discriminative ones (e.g., 64×64). The best overall performance—with an average mIoU of 64.0 / 36.2 / 31.3 on PASCAL VOC, Context, and COCO, respectively—is achieved using the weighted fusion of both cross- and self-attention maps, as indicated by the red-highlighted entry. Finally, attention maps at very low (8×8) or very high (64×64) resolutions individually underperform, reaffirming that mid-level resolutions are the most effective for capturing meaningful object-level features. These findings validate the design choice in SwiftSeg to apply multi-resolution attention fusion, with empirically optimized weights, to maximize segmentation accuracy across diverse datasets.
@@ 4.2.4 Effect of different attention scales
To assess the impact of image description quality on SwiftSeg’s segmentation performance, we experiment with several text prompt generation strategies for constructing the reconstruction prompt $\mathcal{P}$. In addition to a simple text prompt—“A photo of {class candidates}”—which does not rely on any image captioning model, we evaluate a variety of automatic captioning methods. These include BLIP [47], used as the default captioning model in our framework, as well as more recent models such as BLIP-2 [48], GIT [49], ViT-GPT2 [50], and Gemini Flash 2.0 [51]. For Gemini, we consider two prompting strategies: Gemini(1) uses the instruction “Describe this image in one sentence” to generate a full description, whereas Gemini(2) uses “Describe the image in a short and generic sentence, avoid specific names or proper nouns” to generate more abstract, less detailed captions. As shown in Table 7, the model exhibits relatively stable performance across different captioning methods, indicating that SwiftSeg is not overly sensitive to variations in prompt content. Interestingly, we observe that overly detailed prompts (Gemini(1)) slightly degrade performance, whereas more simple prompts (Gemini(2)) yield more better results. Additionally, the simple text prompt also yields suboptimal performance, indicating that purely class-driven prompts without image-specific context are less effective in guiding segmentation. Overall, BLIP provides a balanced captioning quality that is neither too generic nor overly specific, making it a suitable choice for guiding attention in our SwiftSeg.
Table 7: Performance of SwiftSeg on the PASCAL VOC dataset using different methods for generating the text prompt $\mathcal{P}$.
@@
5 Discussion and Conclusion
5.1 Limitations
While SwiftSeg achieves strong segmentation performance and inference efficiency, it still faces several limitations. First, the method is sensitive to the quality of the candidate class list, which must be provided or predicted in advance; incomplete or noisy class sets can lead to false negatives or misclassifications. Second, SwiftSeg relies on static prompt encoding during inference—once the textual prompt is generated, it is not updated or refined adaptively, which may limit its flexibility in complex scenes. Third, as a training-free method, SwiftSeg’s performance is inherently constrained by the representational capacity of the pretrained diffusion backbone (e.g., Stable Diffusion); any limitations or biases in the underlying model directly affect segmentation quality and generalization.
In addition, we present several failure cases of SwiftSeg in Figure 10, which highlight the current limitations of our approach in challenging segmentation scenarios. The first three rows in the left column illustrate difficulties in accurately segmenting object shapes and boundaries, particularly in scenes containing multiple overlapping objects (e.g., the first two rows). These cases suggest limitations in spatial discrimination when semantic boundaries are ambiguous or tightly packed. The top row of the right column demonstrates SwiftSeg’s difficulty in detecting small objects, which may be overlooked due to their weak or missing representation in attention maps. Similarly, the second row of the right column shows segmentation errors when visually similar categories appear in close proximity, indicating challenges in category disambiguation under fine-grained visual similarity. Finally, the bottom row in both columns showcases cases where the quality of the candidate class list heavily influences segmentation performance. In these examples, the failure to predict or include correct class labels in the class prompt results in incomplete or incorrect segmentation, reaffirming SwiftSeg’s dependency on accurate open-set label generation.
Figure 10: Visualization of some failure cases by SwiftSeg

### v3

@@ 4.0.1 Datasets and Metric
We evaluate the proposed FA-Seg on standard benchmark datasets for the Open-Vocabulary Semantic Segmentation (OVSS) task: PASCAL VOC, PASCAL Context, and MS COCO Object, comprising 20, 59, and 80 object categories, respectively (excluding the background class). As FA-Seg is a training-free method, evaluation is conducted directly on the validation sets of each dataset, which include 1,449 images from PASCAL VOC, 5,004 images from PASCAL Context, and 5,000 images from MS COCO Object. To assess segmentation performance and enable fair comparison with existing approaches, we employ the mean Intersection-over-Union (mIoU) metric, which is a widely used standard for evaluating semantic segmentation models that measure the overlap between predicted and ground-truth regions across classes.
@@ 4.0.2 Implementation Details
We employ 2-Rectified Flow [43], a distilled variant of Stable Diffusion 1.5 (SD 1.5), as the primary model for attention map extraction. All inferences of SD are performed at an image resolution of 512$\times$512 pixels. In the multi-resolution attention fusion, we assign distinct weights to cross-attention and self-attention maps. Specifically, for cross-attention maps, we use the weight vector $w^{r}_{\text{cross}}\in\{0.15,0.7,0.15,0\}$ corresponding to resolution levels $r\in\{8,16,32,64\}$, while self-attention maps, we assign weights of $w^{r}_{\text{self}}\in\{0.1,0.1,0.5,0.3\}$. These weights are determined through both qualitative inspection of attention heatmaps and quantitative evaluation of segmentation performance at each resolution (see Table 6). Rather than relying on extensive hyperparameter tuning, we prioritize semantic relevance, assigning greater weights to layers that contribute more effectively to accurate segmentation. All experiments are implemented in PyTorch with a batch size of 2 and conducted on a system equipped with an NVIDIA GeForce RTX 4090 GPU (32GB RAM)
@@ 4.1.1 Quantitative results
Table 1 summarizes the performance of our proposed FA-Seg method across three widely-used open-vocabulary semantic segmentation benchmarks: PASCAL VOC 2012, PASCAL Context, and COCO Object. FA-Seg achieves mIoU scores of 64.0% on PASCAL VOC, 31.3% on PASCAL Context, and 36.2% on COCO Object, leading to the highest average mIoU among training-free diffusion-based approaches across all three datasets. While results from contrastive vision-language models are also reported for comprehensive comparison, our primary focus centers on a direct and fair comparison with recent training-free diffusion-based methods, including Diffusion Baseline [39], FreeSeg-Diff [19], DiffCut [38], DiffSeg [36], DiffSegmenter [20], and InvSeg [21]. The performance of Diffusion Baseline is referenced from the InvSeg study [21], while the results of Semantic DiffSeg, an adaptation of DiffSeg [36] for open-vocabulary settings, are taken from its corresponding publication [20]. FA-Seg establishes state-of-the-art accuracy on PASCAL VOC 2012 and delivers the highest overall average performance across all datasets. On PASCAL Context and COCO Object, FA-Seg ranks second, with results only marginally trailing behind the best-performing methods on each dataset. These strong results stem from our proposed methods: the dual-prompt mechanism improves class-aware attention control, the HARD module enhances object boundary accuracy through multi-resolution attention fusion, and the TTF strategy further improves spatial consistency during inference.
@@
FA-Seg (Ours) RTX 4090 13.4G 0.36s 43.8
Table 3: Ablation study on the contributions of each component in FA-Seg, including self-attention refinement, cross-attention weighted aggregation, self-attention weighted aggregation, class prompt design, and test-time flipping. Removing any individual component leads to decreased performance.
FA-Seg PASCAL VOC PASCAL Context COCO Object $\downarrow\Delta$
@@
w/o cross attention weighted aggregation 61.2 30.2 35.0 1.7
@@
Table 2 presents a quantitative evaluation of the operational efficiency (exclude candidate class extraction) among existing state-of-the-art diffusion-based segmentation methods, including DiffSegmentor, InvSeg, and our FA-Seg . In terms of memory consumption, FA-Seg demonstrates superior efficiency, utilizing only 13.4GB, significantly less than DiffSegmentor’s 16.1GB and InvSeg’s 32.4GB . Notably, regarding inference time, FA-Seg achieves superior performance with just 0.36 seconds on an RTX 4090 GPU. This figure represents a substantial improvement over DiffSegmentor, whose inference time depends on the number of candidate classes M, and is particularly impressive when compared to InvSeg, which requires up to 7.9 seconds per image on an NVIDIA H100 GPU due to its complex prompt embedding optimization procedure. This efficiency stems from FA-Seg’s ability to perform fast semantic reconstruction for all candidate classes in a single inference run. Furthermore, to maximize throughput and scalability, all evaluations are conducted using half-precision inference. While this introduces a marginal trade-off in numerical precision, it results in considerable gains in speed and memory usage. These results demonstrate that FA-Seg offers high inference efficiency while significantly reducing the necessary computational resources.
In summary, FA-Seg not only achieves state-of-the-art segmentation performance among training-free diffusion-based methods but also demonstrates superior inference efficiency and deployment simplicity, making it an efficient and scalable solution for real-world open-vocabulary semantic segmentation.
Figure 8: Open vocabulary semantic segmentation results of FA-Seg and DiffSegmenter [20], DiffCut [38], ClearCLIP [5], CLIP-DIY [3], OVSegmentor [12]
@@ 4.1.2 Qualitative results
Figure 8 provides a qualitative comparison of FA-Seg against several methods, including DiffSegmenter, DiffCut, ClearCLIP, CLIP-DIY, and OVSegmentor, on two benchmark datasets: PASCAL VOC 2012 (top 7 rows) and PASCAL Context (bottom 6 rows). To ensure fairness, the PAMR post-processing step [46] was removed when showing DiffCut results. The results highlight several key observations. DiffSegmenter performs well in capturing object shapes on PASCAL VOC but struggles with contextual classes (e.g., tree, glass) in PASCAL Context. Additionally, its performance is highly sensitive to the quality of the candidate class list. DiffCut produces unstable segmentation masks, attributed to its reliance on an unsupervised grouping strategy and requiring FC-CLIP for semantic label association, making it suboptimal for open-vocabulary segmentation tasks. Among CLIP-based methods, CLIP-DIY often suffers from false positives and false negatives, while ClearCLIP partially mitigates this issue but still shows failure cases—particularly missing detections for classes like person. This indicates limited spatial localization capacity in these models. OVSegmentor exhibits strong geometric segmentation but lacks effective mechanisms to filter out irrelevant or negative classes, leading to noisy outputs and reduced precision.
In contrast, FA-Seg consistently delivers high-quality segmentation masks, demonstrating both accurate geometric localization and precise category alignment, even in challenging scenarios. These results substantiate FA-Seg’s advantage in open-vocabulary settings, offering a strong balance between semantic accuracy and spatial precision.
@@ 4.2.1 Effects of different components
Table  3 presents an ablation study that evaluates the contribution of components in FA-Seg. Removing the cross-attention and self-attention weighted aggregation modules is equivalent to using only a single-resolution attention map with the best performance (as shown in Table  6), without applying multi-resolution aggregation-specifically, $16\times 16$ for cross-attention and $32\times 32\times 32\times 32$ for self-attention. In contrast, removing the class prompt $\mathcal{P}^{\prime}$ means that cross-attention maps are computed solely based on the text prompt $\mathcal{P}$, without explicitly guiding the model to focus on candidate classes. In the original configuration, the class candidate list is appended to the image caption to form the complete prompt $\mathcal{P}$, allowing the model to utilize the appended candidate list to generate more targeted cross-attention maps. Among all components, self-attention refinement proves to be the most critical, as its removal leads to the most significant average performance drop (-15.1%), underscoring its importance in enhancing the spatial precision of class-aware attention. Both cross-attention and self-attention weighted aggregation also play important roles, as their removal results in noticeable performance degradation. On the other hand, class prompt design and test-time flipping (TTF) have relatively more minor effects-0.6% and 0.3%, respectively-yet consistently contribute to overall segmentation quality.
These findings validate the overall design of FA-Seg, where each component serves a complementary role in improving segmentation performance while maintaining computational efficiency.
4.2.2 Different background threshold
@@
4.2.3 Different class candidate generation strategies.
Table 5: Ablating different class candidate generation strategies.
PASCAL VOC VOC Context COCO Object
BIP 58.3 26.4 34.3
CLIP 63.0 23.9 34.8
TagCLIP 65.0 29.3 34.8
CLIP+BIP 62.8 28.8 36.7
TagCLIP+BIP 64.0 31.3 36.2
Ground Truth 69.7 45.1 38.8
4.2.4 Effect of different attention scales
Table 6 presents a comprehensive study on the effect of different attention scales and fusion strategies in FA-Seg. The results reveal several key findings. First, single-resolution attention maps, particularly those from the 16×16 resolution level, consistently yield higher segmentation performance across all datasets compared to other individual scales, suggesting that this resolution offers a favorable balance between semantic context and spatial detail. Second, when combining attention maps across multiple resolutions, both mean and weighted aggregation strategies improve overall performance, highlighting the complementary nature of multi-scale attention. Notably, weighted fusion outperforms mean fusion, as it enables the model to emphasize semantically rich and informative layers (e.g., 16×16 and 32×32) while down-weighting noisier or less discriminative ones (e.g., 64×64). The best overall performance—with an average mIoU of 64.0 / 36.2 / 31.3 on PASCAL VOC, Context, and COCO, respectively—is achieved using the weighted fusion of both cross- and self-attention maps, as indicated by the red-highlighted entry. Finally, attention maps at very low (8×8) or very high (64×64) resolutions individually underperform, reaffirming that mid-level resolutions are the most effective for capturing meaningful object-level features. These findings validate the design choice in FA-Seg to apply multi-resolution attention fusion, with empirically optimized weights, to maximize segmentation accuracy across diverse datasets.
@@ 4.2.4 Effect of different attention scales
To assess the impact of image description quality on FA-Seg’s segmentation performance, we experiment with several text prompt generation strategies for constructing the reconstruction prompt $\mathcal{P}$. In addition to a simple text prompt—“A photo of {class candidates}”—which does not rely on any image captioning model, we evaluate a variety of automatic captioning methods. These include BLIP [47], used as the default captioning model in our framework, as well as more recent models such as BLIP-2 [48], GIT [49], ViT-GPT2 [50], and Gemini Flash 2.0 [51]. For Gemini, we consider two prompting strategies: Gemini(1) uses the instruction “Describe this image in one sentence” to generate a full description, whereas Gemini(2) uses “Describe the image in a short and generic sentence, avoid specific names or proper nouns” to generate more abstract, less detailed captions. As shown in Table 7, the model exhibits relatively stable performance across different captioning methods, indicating that FA-Seg is not overly sensitive to variations in prompt content. Interestingly, we observe that overly detailed prompts (Gemini(1)) slightly degrade performance, whereas more simple prompts (Gemini(2)) yield more better results. Additionally, the simple text prompt also yields suboptimal performance, indicating that purely class-driven prompts without image-specific context are less effective in guiding segmentation. Overall, BLIP provides a balanced captioning quality that is neither too generic nor overly specific, making it a suitable choice for guiding attention in our FA-Seg.
5 Discussion and Conclusion
5.1 Limitations
While FA-Seg achieves strong segmentation performance and inference efficiency, it still faces several limitations. First, the method is sensitive to the quality of the candidate class list, which must be provided or predicted in advance; incomplete or noisy class sets can lead to false negatives or misclassifications. Second, FA-Seg relies on static prompt encoding during inference—once the textual prompt is generated, it is not updated or refined adaptively, which may limit its flexibility in complex scenes. Third, as a training-free method, FA-Seg’s performance is inherently constrained by the representational capacity of the pretrained diffusion backbone (e.g., Stable Diffusion); any limitations or biases in the underlying model directly affect segmentation quality and generalization.
Table 7: Performance of FA-Seg on the PASCAL VOC dataset using different methods for generating the text prompt $\mathcal{P}$.
@@ 5.1 Limitations
In addition, we present several failure cases of FA-Seg in Figure 10, which highlight the current limitations of our approach in challenging segmentation scenarios. The first three rows in the left column illustrate difficulties in accurately segmenting object shapes and boundaries, particularly in scenes containing multiple overlapping objects (e.g., the first two rows). These cases suggest limitations in spatial discrimination when semantic boundaries are ambiguous or tightly packed. The top row of the right column demonstrates FA-Seg’s difficulty in detecting small objects, which may be overlooked due to their weak or missing representation in attention maps. Similarly, the second row of the right column shows segmentation errors when visually similar categories appear in close proximity, indicating challenges in category disambiguation under fine-grained visual similarity. Finally, the bottom row in both columns showcases cases where the quality of the candidate class list heavily influences segmentation performance. In these examples, the failure to predict or include correct class labels in the class prompt results in incomplete or incorrect segmentation, reaffirming FA-Seg’s dependency on accurate open-set label generation.